# Data Mixing Agent: Learning to Re-weight Domains for Continual Pre-training

### v1

@@ Abstract
Continual pre-training on small-scale task-specific data is an effective method for improving large language models in new target fields, yet it risks catastrophic forgetting of their original capabilities. A common solution is to re-weight training data mixtures from source and target fields on a domain space to achieve balanced performance. Previous domain reweighting strategies rely on manual designation with certain heuristics based on human intuition or empirical results. In this work, we prove that more general heuristics can be parameterized by proposing Data Mixing Agent, the first model-based, end-to-end framework that learns to re-weight domains. The agent learns generalizable heuristics through reinforcement learning on large quantities of data mixing trajectories with corresponding feedback from an evaluation environment. Experiments in continual pre-training on math reasoning show that Data Mixing Agent outperforms strong baselines in achieving balanced performance across source and target field benchmarks. Furthermore, it generalizes well across unseen source fields, target models, and domain spaces without retraining. Direct application to the code generation field also indicates its adaptability across target domains. Further analysis showcases the agents’ well-aligned heuristics with human intuitions and their efficiency in achieving superior model performance with less source-field data.
@@ 1 Introduction
Modern Large Language Models (LLMs) are usually pre-trained with trillion-token large-scale general domain datasets (Yang et al., 2025; Liu et al., 2024a). Despite with strong generalization capabilities, these foundation models often require further enhancement in certain knowledge-intensive fields, such as applications in math problem solving (Yang et al., 2024; Shao et al., 2024) or code generation (Guo et al., 2024; Hui et al., 2024). Due to the formidable cost of pre-training from scratch, adapting foundation models to new knowledge/capabilities is usually achieved via continual pre-training on smaller-scale, high-quality data in the target field. However, directly adapting to the target field data can lead to catastrophic forgetting (Dyer et al., 2022) of source data and collapse on existing model capabilities (Lin et al., 2025), normally due to the significant distribution shift between source and target fields.
@@
A popular solution is to curate data mixtures of the source and target fields to achieve a balanced performance (Shi et al., 2024). Existing methods mainly organize data mixtures defined by meta-attributes such as data sources and focus, known as domains (Du et al., 2022; Luo et al., 2024). During training, the data mixture is allocated through a distribution in the domain space, which reflects the ratio of data allocated from each domain by reweighting. The distribution can be adjusted after several training steps when necessary, leading to a data mixing trajectory (Luo et al., 2024; Xia et al., 2023) along the domain reweighting steps. Data mixing trajectories are proven to significantly influence model performance (OLMo et al., 2024; Grattafiori et al., 2024; Li et al., 2024), which leads previous works to explore various data mixing algorithms for determining the optimal trajectory for various tasks (Liu et al., 2024b; Ye et al., 2024; Xie et al., 2023; Luo et al., 2024; Xia et al., 2023).
A commonality of these data mixing methods is that their designations are all based on certain general heuristics, such as "data mixtures that provide balanced evaluation loss lead to desired downstream performance". Another indication of these heuristics is the various empirical conclusions drawn from training practices. For example, Wettig et al. (2025) concluded that "Data from Science domain heavily promote model performance on MMLU (Hendrycks et al., 2020), while the Home domain improves HellaSwag (Zellers et al., 2019) performance". In Fig. 1, we further provide an average of distributions along 20 randomly generated data mixing trajectories, separated by increasing/decreasing performance on the MMLU and MATH (Hendrycks et al., 2021) benchmarks. The results show an explicit gap between data distributions that increase/decrease model performance. For example, in MMLU, increasing the DCLM data from the Science and Home&\&&Garden domains improves benchmark performance. In MATH, increasing Dolmino-math data from the Hobbies&\&&Leisure and Real estate domains while keeping a balanced mix of DCLM data is likely to improve benchmark performance. These results prove the existence of more general heuristics for data mixing in continual pre-training.
@@
Figure 1: Four averaged distributions drawn from 20 randomly generated data mixing trajectories. Each distribution in the trajectories is first categorized by whether it increases/decreases the performance of a 50M target model on the MMLU/MATH benchmarks within one re-weighting step. Each category is then averaged to obtain the corresponding distribution in the figure. The models are trained on a 52-dimensional space (more details in Sec. 3.1), mixing the DCLM (Li et al., 2024) and the math split of the Dolmino-mix-1124 (OLMo et al., 2024) dataset.
@@
The above arguments reveal a potentially rich heuristic space for domain reweighting. We believe these model- and data-agnostic heuristics can be unified into a small agent model to guide the data mixing trajectories in an end-to-end manner. Based on this intuition, we propose Data Mixing Agent, the first model-based method that learns to re-weight domains for continual pre-training. We start by randomly sampling large quantities of data mixing trajectories, each with fixed domain re-weighting steps. We then train small proxy models on all trajectories, obtaining model checkpoints on each re-weighting step. All checkpoints are then evaluated on a light-weight but accurate evaluation environment to assess targeted capabilities. The sampled trajectories and corresponding environment feedback are expected to empirically enclose a considerable range of heuristics. Data Mixing Agent is then trained via random sampling on these collected data, and optimized in an off-policy reinforcement learning manner using the Conservative Q-Learning (CQL) algorithm (Kumar et al., 2020). During continual pre-training on the target model, Data Mixing Agent directly predicts the domain distribution for the next data reweighting step on the fly, considering previous states in the data mixing trajectory and the environment feedback.
@@
We apply Data Mixing Agent for continual pre-training on the math reasoning target field while preserving performance in the general field. Evaluation on in-distribution source field data shows that Data Mixing Agent significantly outperforms the strong RegMix (Liu et al., 2024b) baseline, achieving an average improvement of 3.02% across 8 general benchmarks and 4 math reasoning benchmarks. The agent’s generalization ability is demonstrated by its balanced performance across 3 unseen source fields, 4 target models, and 2 domain spaces, all without retraining. We further use the trained agent on the math reasoning field directly to the unseen code generation target field, showing that its learned heuristics partially generalize across target domains. Additional analysis confirms that these heuristics align well with human intuitions, and Data Mixing Agent can efficiently leverage the data mixture to achieve superior continual pre-training performance with less source-field data.
In summary, this work makes the following contributions:
@@
  • We propose the framework of Data Mixing Agent, the first model-based, lightweight domain reweighting method for continual pre-training, guiding the training recipe for the target model in an end-to-end manner;
  • Extensive experiments prove the effectiveness of data mixing agents in alleviating catastrophic forgetting in continual pre-training and achieving balanced performance across model capabilities;
@@
  • Data mixing agent learns heuristics that generalize across source and target fields, target models, and domain spaces, enabling its application in multiple scenarios.
2 Domain Re-weighting as Markov Decision Process
@@
The state space $\mathcal{S}$ is a continuous space consisting of all data distributions from previous domain reweighting steps. Specifically, for step $t$, the state $s_{t}\in\mathcal{S}$ has dimension $s_{t}\in\mathcal{R}^{t\times N}$, where $N$ is the dimension of the action space, determined by the definition of the domains.
@@
The action space $\mathcal{A}$ is a continuous space denoting the data distribution in the current domain reweighting step. At step $t$, the action $a_{t}\in\mathcal{A}$ is a probability distribution over the domain space: $a_{t}\in\mathcal{R}^{N}$ and $\sum_{i=1}^{N}a_{t}^{i}=1$, $a_{t}^{i}\geq 0$.
@@
The policy function $f$ denotes the model-based agent for guiding domain reweighting, which directly determines the action at the current step based on previous states and environment feedback. The feedback are modeled by the reward function $r$, determined by the target fields for continual pre-training and the evaluation environment design. With domain re-weighting modeled as a MDP, the data mixing agent can be optimized via reinforcement learning (Kumar et al., 2020).
@@
The start state $\rho_{s}$ depends on the target model checkpoint, reflecting the domain distribution of its pre-training data. The terminate state $\rho_{s}$ depends on manual setting or data scale, as the training process can often end with predefined token budgets or exhaustion of target field data.
@@ 3 Data Mixing Agent
In this section, we introduce the methodology of Data Mixing Agent, mainly including three procedures: 1) modeling the heuristic space by randomly sampling data mixing trajectories and collecting feedback from an evaluation environment; 2) parameterizing the heuristic space by training the model-based agent on the collected trajectories and feedback via conservative Q-learning; 3) utilizing the data mixing agent to guide the domain reweighting for the target models on the fly to achieve balanced performance. An overview of the pipeline is shown in Fig. 2.
Figure 2: An overview of the training and domain reweighting pipeline of the data mixing agent. We first sample large quantities of data mixing trajectories and train small proxy models on them. Each model checkpoint obtains feedback from the evaluation environment. Secondly, the data mixing agent is optimized on these trajectories and feedback via supervised fine-tuning and off-policy reinforcement learning with a CQL-based Q function. During guiding the training of the target model, the agent directly determines the distribution for the next domain re-weighting step on the fly.
3.1 Modeling the Heuristic Space with Trajectory Sampling

### v2

@@ Abstract
Continual pre-training on small-scale task-specific data is an effective method for improving large language models in new target fields, yet it risks catastrophic forgetting of their original capabilities. A common solution is to re-weight training data mixtures from source and target fields on a domain space to achieve balanced performance. Previous domain reweighting strategies rely on manual designation with certain heuristics based on human intuition or empirical results. In this work, we prove that more general heuristics can be parameterized by proposing Data Mixing Agent, the first model-based, end-to-end framework that learns to re-weight domains. The agent learns generalizable heuristics through reinforcement learning on large quantities of data mixing trajectories with corresponding feedback from an evaluation environment. Experiments in continual pre-training on math reasoning show that Data Mixing Agent outperforms strong baselines in achieving balanced performance across source and target field benchmarks. Furthermore, it generalizes well across unseen source fields, target models, and domain spaces without retraining. Direct application to the code generation field also indicates its adaptability across target domains. Further analysis shows the agents’ well-aligned heuristics with human intuitions and efficiency in achieving superior model performance with less source-field data.
Data Mixing Agent: Learning to Re-weight Domains for Continual Pre-training
Kailai Yang1*  Xiao Liu2   Lei Ji2  Hao Li3   Xiao Liang4  Zhiwei Liu1 Yeyun Gong2Peng Cheng2Mao Yang2 1 The University of Manchester 2 Microsoft Research 3 Imperial College London 4 University of California, Los Angeles
11footnotetext: Work done during his internship at Microsoft Research.22footnotetext: Corresponding authors.
@@ 1 Introduction
Large Language Models (LLMs) (Yang et al., 2025; Liu et al., 2024a), though pre-trained to obtain generalization capabilities, often require further enhancement in knowledge-intensive fields (Yang et al., 2024; Guo et al., 2024) via continual pre-training in the target field. However, directly adapting to the target field data can lead to catastrophic forgetting of source data and collapse on existing model capabilities (Hui et al., 2024; Lin et al., 2025), due to the significant distribution shift between source and target fields.
@@
A popular solution is to curate data mixtures of the source and target fields to achieve a balanced performance (Shi et al., 2024). Existing methods mainly organize data mixtures defined by meta-attributes such as data sources and focus, known as domains (Du et al., 2022; Luo et al., 2024). During training, the data mixture is allocated through a distribution in the domain space, which reflects the ratio of data allocated from each domain. The distribution can be adjusted after several training steps if necessary, leading to a data mixing trajectory (Luo et al., 2024; Xia et al., 2023) along the domain reweighting steps. Data mixing trajectories significantly influence model performance (OLMo et al., 2024; Grattafiori et al., 2024; Li et al., 2024), and previous works have explored data mixing algorithms to determine optimal trajectories for different tasks (Liu et al., 2024b; Ye et al., 2024; Xie et al., 2023; Xia et al., 2023; Luo et al., 2024).
@@
Figure 1: Four averaged distributions drawn from 20 randomly generated data mixing trajectories. Each distribution in the trajectories is first categorized by whether it increases/decreases the performance of a 50M target model on the MMLU or MATH benchmarks within one re-weighting step, leading to four categories. The distributions in each category are then averaged within the domain space to represent their features, as shown in the figure. The models are trained on a 52-dimensional space, mixing the general domain data from DCLM (Li et al., 2024) and the math reasoning data from the math split of the Dolmino-mix-1124 (OLMo et al., 2024) dataset.
A commonality of these data mixing methods is that their designs are based on certain general heuristics, such as: data mixtures that provide balanced evaluation loss lead to desired downstream performance. Another indication of these heuristics is the various empirical conclusions drawn from training practices. For example, Wettig et al. (2025) concluded that Data from the ‘Science’ domain heavily promote model performance on MMLU (Hendrycks et al., 2020), while the ‘Home’ domain improves HellaSwag (Zellers et al., 2019) performance. In Fig. 1, we provide an average of distributions along 20 randomly generated data mixing trajectories (each with 80 domain reweighting steps), separated into four categories. The categories are defined by whether they increase or decrease the target model’s performance on the MMLU/MATH (Hendrycks et al., 2021) benchmarks. The data mixing domains are defined and classified by the Nvidia domain classifier  [^1] https://huggingface.co/nvidia/domain-classifier. According to the results, there are explicit differences between data distributions that increase/decrease model performance. For example, in MMLU (the first and second rows), higher ratios of DCLM data from the Science and Home&\&Garden domains significantly improves benchmark performance. In MATH, increasing data mixtures from the Hobbies&\&Leisure and Real estate domains of the Dolmino-math data while keeping a balanced mix of DCLM data is likely to boost benchmark performance. The above results unveil more general heuristics that can enhance model performance in continual pretraining, yet have not been discovered or utilized by existing works. This negligence motivates further efforts to efficiently mine such heuristics and leverage them for data mixing in different scenarios.
@@
These observations reveal a rich heuristic space for domain reweighting. We believe these model- and data-agnostic heuristics can be unified into a small agent model to guide the data mixing trajectories in an end-to-end manner. Based on this intuition, we propose Data Mixing Agent, the first model-based method that learns to re-weight domains for continual pre-training. We start by randomly sampling large quantities of data mixing trajectories, each with fixed domain re-weighting steps. We then train small proxy models on all trajectories, obtaining model checkpoints on each re-weighting step. All checkpoints are evaluated in a light-weight evaluation environment to assess the target capabilities. The trajectories and corresponding environment feedback are expected to empirically enclose a wide range of heuristics. Data Mixing Agent is then optimized on these collected data with the Conservative Q-Learning (CQL) reinforcement learning algorithm (Kumar et al., 2020). During continual pre-training on the target model, the agent directly predicts the domain distribution for the next domain reweighting step on the fly, considering previous states in the data mixing trajectory and the environment feedback.
@@
We apply Data Mixing Agent on the math reasoning target field while preserving performance in the general field. Evaluation on in-distribution source field data shows that our method significantly outperforms state-of-the-art static (Liu et al., 2024b) and dynamic (Xia et al., 2023) domain reweighting methods, achieving an average improvement of 3.02% across 8 general benchmarks and 4 math reasoning benchmarks. The agent’s generalization ability is demonstrated with balanced performance across 3 unseen source fields, 4 target models, and 2 domain spaces, all without retraining. We also directly apply agents trained on math reasoning to guide training on the unseen code generation field, proving their generalization across target domains. Additional analysis confirms that these heuristics align well with human intuitions, and Data Mixing Agent can achieve superior continual pre-training performance with less source-field data.
In summary, we make the following contributions:
@@
  • We propose Data Mixing Agent, the first end-to-end, lightweight domain reweighting method for continual pre-training;
  • Data Mixing Agent significantly outperforms state-of-the-art baselines in extensive experiments, leading to balanced performance across model capabilities and high efficiency in data usage;
@@
  • Data Mixing Agent learns heuristics that generalize across source and target fields, target models, and domain spaces.
2 Domain Re-weighting as MDP
@@
The state space $\mathcal{S}$ is a continuous space consisting of all data distributions from previous domain reweighting steps. Specifically, for step $t$, the state $s_{t}\in\mathcal{S}$ has dimension $s_{t}\in\mathbb{R}^{t\times N}$, where $N$ is the dimension of the action space, determined by the definition of the domains.
@@
The action space $\mathcal{A}$ is a continuous space denoting the data distribution in the current domain reweighting step. At step $t$, the action $a_{t}\in\mathcal{A}$ is a probability distribution over the domain space: $a_{t}\in\mathbb{R}^{N}$ and $\sum_{i=1}^{N}a_{t}^{i}=1$, $a_{t}^{i}\geq 0$.
@@
The policy function $f$ denotes Data Mixing Agent that determines the action at the current step. The feedback is modeled by the reward function $r$, determined by the target fields and the environment. With domain re-weighting modeled as an MDP, the policy function can be optimized via reinforcement learning.
@@
The start state $\rho_{s}$ depends on the target model, reflecting the distribution of its pre-training data. The terminate state $\rho_{e}$ depends on manual setting or data scale, as the training process can often end with predefined token budgets or the exhaustion of target field data.
@@ 3 Data Mixing Agent
In this section, we introduce the methodology of Data Mixing Agent. An overview of the pipeline is shown in Fig. 2. The full implementation details of all procedures are described in Appn. C.
Figure 2: An overview of Data Mixing Agent. We first sample data mixing trajectories and train small proxy models on them. Each model checkpoint obtains feedback from the evaluation environment. Secondly, the agent is optimized on these trajectories and feedback via CQL. When guiding continual pretraining, the agent determines the distribution for the next step on the fly.
3.1 Trajectory Sampling